# Fully stabilized Minkowski vacua in the $2^{6}$ Landau-Ginzburg model

### v1

@@ 5 Conclusion
In this paper, we have studied flux compactifications on an orientifold of the $2^{6}$ Landau-Ginzburg model. We focused on 4d $\mathcal{N}=1$ Minkowski vacua that arise for $G_{3}$ fluxes that are imaginary-self-dual. We demonstrated that, for several flux choices, there is non-trivial moduli stabilization from higher-than-quadratic order terms in the expansion of the superpotential around a critical point. This means that there are massless fields that are stabilized by higher order interaction terms, similar to a massless $\phi^{4}$ scalar field that has no flat direction. Unlike in the $1^{9}$ Landau-Ginzburg model Becker:2024ijy , it is not tractable to study all flux examples beyond terms in the superpotential that are cubic in the complex structure moduli. However, we find two interesting solutions in the $2^{6}$ model that violate the Refined Tadpole Conjecture. One example has 52 massive fields for a flux choice that only contributes $N_{\rm flux}=16$ to the O3 tadpole condition. Another example has $N_{\rm flux}=28$, and 75 massive fields but has at least $84=3\cdot 28$ stabilized fields, which again violates the Refined Tadpole Conjecture. Nevertheless, it seems that this model is actually confirming the original Tadpole Conjecture, with a modified bound, as is shown by studying a large number of flux configurations in Becker2024 .
@@ Acknowledgments
We like to thank Katrin Becker, Mariana Graña, Miguel Montero, Miguel Morros, Sav Sethi, Johannes Walcher, and Qi You for useful discussions. The work of AS is supported in part by the NSF grant PHY-2112859. MR acknowledges the support of the Dr. Hyo Sang Lee Graduate Fellowship from the College of Arts and Sciences at Lehigh University. The work of MR and TW is supported in part by the NSF grant PHY-2210271. MR and TW would like to thank the Erwin-Schrödinger-Institute in Vienna for their hospitality; this work was completed there during the program “The Landscape vs. the Swampland”.

### v2

@@ 5 Conclusion
In this paper, we have studied flux compactifications on an orientifold of the $2^{6}$ Landau-Ginzburg model. We focused on 4d $\mathcal{N}=1$ Minkowski vacua that arise for $G_{3}$ fluxes that are imaginary-self-dual. We demonstrated that, for several flux choices, there is non-trivial moduli stabilization from higher-than-quadratic order terms in the expansion of the superpotential around a critical point. This means that there are massless fields that are stabilized by higher order interaction terms, similar to a massless $\phi^{4}$ scalar field that has no flat direction. Unlike in the $1^{9}$ Landau-Ginzburg model Becker:2024ijy , it is not tractable to study all flux examples beyond terms in the superpotential that are cubic in the complex structure moduli. However, we find two interesting solutions in the $2^{6}$ model that violate the Refined Tadpole Conjecture. One example has 52 massive fields for a flux choice that only contributes $N_{\rm flux}=16$ to the O3 tadpole condition. Another example has $N_{\rm flux}=28$, and 75 massive fields but has at least $84=3\cdot 28$ stabilized fields. Both examples violates the Refined Tadpole Conjecture by more than a factor of 2. Nevertheless, it is possible that this model is actually confirming the original Tadpole Conjecture, with a modified bound, as is shown by studying a large number of flux configurations in Becker:2024ayh .
@@ Acknowledgments
We like to thank Katrin Becker, Mariana Graña, Daniel Junghans, Miguel Montero, Miguel Morros, Sav Sethi, Johannes Walcher, and Qi You for useful discussions. The work of AS is supported in part by the NSF grant PHY-2112859. MR acknowledges the support of the Dr. Hyo Sang Lee Graduate Fellowship from the College of Arts and Sciences at Lehigh University. The work of MR and TW is supported in part by the NSF grant PHY-2210271. MR and TW would like to thank the Erwin-Schrödinger-Institute in Vienna for their hospitality; this work was completed there during the program “The Landscape vs. the Swampland”.